OpenStudy (shealyn):

What is the solution of 2(x - 5) + 4 = 30?

OpenStudy (jtug6):

multiply the 2 inside the the parenthesis first. 2*x and 2*(-5). And then add those. What do you get?

OpenStudy (shealyn):

2x+-10??????????/

OpenStudy (jtug6):

Yes! When you add a negative, you can just say the same thing as subtracting. So, 2x-10! Awesome. So, we still have that +4 though so we tag it on too. 2x-10+4. What's -10+4?

OpenStudy (shealyn):

-6

OpenStudy (jtug6):

yup! so now it's just, 2x-6 = 30. So we need to solve for x. What do you think we should do to both sides first?

OpenStudy (shealyn):

add 6

OpenStudy (jtug6):

Yes! Great. So, 2x-6+6 = 30 + 6, right? so, 2x = 36. Now what? ^^

OpenStudy (shealyn):

divide 2 on both sides

OpenStudy (jtug6):

yes! . x = 36/2 = ?

OpenStudy (shealyn):

x=6

OpenStudy (jtug6):

mmm, that'd be if it was 36/6, but it's 36/2 actually :P

OpenStudy (shealyn):

WAIT DUH ITS 18

OpenStudy (jtug6):

mhm! x = 18 :D

OpenStudy (shealyn):

Tysm lmao

OpenStudy (jtug6):

The great thing is, you can test it. Plug x = 18 back into the very first equation ;P

OpenStudy (jtug6):

See what you get. 2(18-5) + 4 = 30?

OpenStudy (jtug6):

2(13) + 4 = 30 26 + 4 = 30 ~ 30 = 30 ~

OpenStudy (shealyn):

thank you

OpenStudy (jtug6):

no problem!